Causwell Company began 2021 with 26,000 units of inventory on hand. The cost of each unit was $5.00. During 2021 an additional 46,000 units were purchased at a single unit cost, and 36,000 units remained on hand at the end of 2021 (36,000 units therefore were sold during 2021). Causwell uses a periodic inventory system. Cost of goods sold for 2021, applying the average cost method, is $221,400. The company is interested in determining what cost of goods sold would have been if the FIFO or LIFO methods were used.

Required: 1. Determine the cost of goods sold for 2021 using the FIFO method. [Hint: Determine the cost per unit of 2021 purchases.] (Do not round intermediate calculations.)
Answer #1

Correct Answer:

Cost of Goods Sold (FIFO)

$ 198,000.00

Working:

A

cost of goods sold using weighted average method

$     221,400.00

B

total units sold

36000

C =A/B

cost per unit

$                  6.15

D

total units available for sale

72000

E=C*D

total cost of goods available for sale

$     442,800.00

F

Less: cost of beginning inventory

$     130,000.00

G =E-F

Total cost of46000 units purchased

$     312,800.00

H = G/46000

per unit cost of 46000 units

$                  6.80

Cost of Goods Available for sale

Units

Cost per unit

value

Beginning Inventory

26000

$                       5.00

$              130,000.00

Purchases

46000

$                       6.80

$              312,800.00

Total

72000

$              442,800.00

FIFO

A

Total Units Available for sale

72000

$ 442,800.00

Units Sold

36000

Ending Inventory Units

36000

Valuation

Cost of Goods Sold

26000

$                           5.00

$ 130,000.00

10000

$                           6.80

$    68,000.00

B

Cost of Goods Sold

36000

units

$ 198,000.00

A-B

Ending Inventory

36000

units

$ 244,800.00

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks